1. y(y*5)(wy)
2. |6-3.7|+|7-(5-2)|
3. 8-3|5-16+1|
This is my first Algebra course. Kindly assist me. Thank you

Answer by stanbon(57387)

1. y(y*5)(wy) = 5y^2(wy) = 5wy^3
2. |6-3.7|+|7-(5-2)| = |2.3] + |7-3| = 2.3 + 4 = 6.3
3. 8-3|5-16+1| = 8-3(10) = -22
================
Cheers,
Stan H.